Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A drawer holds 4 red hats and 4 blue hats. What is the...

Expert replies
by AAPL » Wed Nov 29, 2017 6:51 pm
A drawer holds 4 red hats and 4 blue hats. What is the probability of getting exactly three red hats or exactly three blue hats when taking out 4 hats randomly out of the drawer and returning each hat before taking out the next one?

A. 1/8
B. 1/4
C. 1/2
D. 3/8
E. 7/12

The OA is C.

Please, can any expert assist me with this PS question? I don't have it clear and I appreciate if any explain it for me. Thanks.
Join the discussion
Source: — Problem Solving |

by [email protected] » Wed Nov 29, 2017 7:03 pm
Hi swerve,

We're told that a drawer holds 4 red hats and 4 blue hats. We're asked for the probability of getting exactly three red hats or exactly three blue hats when taking out 4 hats randomly out of the drawer and returning each hat before taking out the next one. We can answer this question with a bit of note-taking and some basic arithmetic. To start, since we're putting each hat back into the drawer after pulling one out, the fact that there are 4 red hats and 4 blue hats is irrelevant. As long as there are an equal number of both colors, the math will be exactly the same (meaning that there could be 1 red hat and 1 blue hat).

Since there is an equal chance of pulling either of the two colors on each 'pull', there are (2)(2)(2)(2) = 16 possible outcomes

Given the 'goal', there are only a limited number of possible outcomes that 'fit' what we're looking for:
RRRB
RRBR
RBRR
BRRR

BBBR
BBRB
BRBB
RBBB

8 of the 16 = 1/2 of the outcomes give us what we're after.

Final Answer: C

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by GMATGuruNY » Thu Nov 30, 2017 4:00 am
AAPL wrote:A drawer holds 4 red hats and 4 blue hats. What is the probability of getting exactly three red hats or exactly three blue hats when taking out 4 hats randomly out of the drawer and returning each hat before taking out the next one?

A. 1/8
B. 1/4
C. 1/2
D. 3/8
E. 7/12
[/quote

P(exactly n times) = P(one way) * total possible ways.

Since 4 of the 8 hats are red, and 4 of the 8 hats are blue:
P(R) = 4/8 = 1/2.
P(B) = 4/8 = 1/2.

P(one way):
One way to select exactly 3 blue hats:
RBBB.
P(R on the 1st pick) = 1/2.
P(B on the 2nd pick) = 1/2.
P(B on the 3rd pick) = 1/2.
P(B on the 4th pick) = 1/2.
Since we want all of these events to happen, we MULTIPLY the fractions:
1/2 * 1/2 * 1/2 * 1/2 = 1/16.

Total possible ways:
RBBB is only ONE WAY to get exactly 3 blue hats.
Now we must account for ALL OF THE WAYS to get exactly 3 blue hats.
Any arrangement of the letters RBBB represents one way to get exactly 3 blue hats.
Thus, to account for ALL OF THE WAYS to get exactly 3 blue hats, the result above must be multiplied by the number of ways to arrange the letters RBBB.
Number of ways to arrange 4 elements = 4!.
But when an arrangement includes IDENTICAL elements, we must divide by the number of ways each set of identical elements can be ARRANGED.
The reason:
When the identical elements swap positions, the arrangement doesn't change.
Here, we must divide by 3! to account for the three identical B's:
4!/(3!) = 4.

Multiplying the results above, we get:
P(exactly 3 blue hats) = 1/16 * 4 = 1/4.

By extension:
P(exactly 3 red hats) = 1/4.

Since a favorable outcome will be yielded by selecting exactly 3 blue hats OR exactly 3 red hats, we ADD the fractions in blue:
P(3 hats of the same color) = P(exactly 3 blue hats) + P(exactly 3 red hats) = 1/4 + 1/4 = 1/2.

The correct answer is C.

More practice:
https://www.beatthegmat.com/select-exact ... 88786.html
https://www.beatthegmat.com/probability- ... 14250.html
https://www.beatthegmat.com/a-single-par ... 28342.html
https://www.beatthegmat.com/at-a-blind-t ... 20058.html
https://www.beatthegmat.com/rain-check-t79099.html
https://www.beatthegmat.com/probability-t227448.html
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Scott@TargetTestPrep » Sat Oct 05, 2019 3:54 pm
AAPL wrote:A drawer holds 4 red hats and 4 blue hats. What is the probability of getting exactly three red hats or exactly three blue hats when taking out 4 hats randomly out of the drawer and returning each hat before taking out the next one?

A. 1/8
B. 1/4
C. 1/2
D. 3/8
E. 7/12

The OA is C.

Please, can any expert assist me with this PS question? I don't have it clear and I appreciate if any explain it for me. Thanks.
First, let’s determine the probability of selecting exactly 3 red hats (and 1 blue hat).

P(R-R-R-B) = (1/2)^4 = 1/16

Since R-R-R-B can be arranged in 4!/3! = 4 ways, the overall probability is 4/16 = 1/4.

Since we have the same number of red and blue hats the probability of selecting exactly 3 blue hats (and 1 red hat) is also 1/4.

Thus, the probability of selecting exactly 3 red hats or 3 blue hats is 1/4 + 1/4 = 2/4 = 1/2.

Answer: C

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion